Ann got a 15 year Fully Amortizing FRM for $1,000,000 at an annual interest rate of 7% compounded monthly, with monthly payments. After 5 years of payments, Ann can refinance the balance into a 10 year Fully Amortizing FRM at an annual interest rate of 5.25% compounded monthly, with monthly payments. If Ann refinances into this 10 year loan, what will be her monthly savings on her mortgage payment? "
Answer #1

USING FINANCIAL CALCULATOR
Step 1: Monthly payment initially
N=12*15
I/Y=7%/12
PV=-1000000
FV=0
CPT PMT=8988.28270852427

Step 2: Loan balance after 1st year
N=12*5
I/Y=7%/12
PV=-1000000
PMT=8988.28270852427
CPT FV=774128.019677181

Step 3: New monthly payment
N=12*10
PV=-774128.019677181
I/Y=5.25%/12
FV=0
CPT PMT=8305.75

USING EXCEL
=-PMT(5.25%/12,12*10,FV(7%/12,12*5,PMT(7%/12,12*15,-1000000),-1000000))
=8305.75